OpenStudy (anonymous):

Match the lengths of the hypotenuse and one leg of a right triangle to the length of the other leg. 1. Hypotenuse: 10; Leg: 6 A. 8 B. 5 C. 7 D. 4

OpenStudy (shamim):

8

OpenStudy (anonymous):

how did you find that answer? @shamim

OpenStudy (shamim):

Can i show steps of my work

OpenStudy (anonymous):

Yes please @shamim

OpenStudy (shamim):

10^2=6^2+a^2

OpenStudy (shamim):

a^2=64

OpenStudy (shamim):

a=8

OpenStudy (shamim):

I used pithagoras formula

OpenStudy (shamim):

Did u get it

OpenStudy (shamim):

Hipoteneus ^2=leg^2+oposit^2